A. Identify the proportion involved in the problem. Write DP for direct proportion, IP for an indirect proportion, and PP for a partitive proportion. Then solve the problem



1. If it takes 3 men to build a fence in 8 hours, how long will it take 6 men to complete the same job?


2. An amount of P5000 is divided between mina and tina in the ratio 3: 5. How much does each girl get?


3. Rolando is a salesman in a department store. He is paid P75 per hour. How long must he work to earn P600 a day?


4. The ratio of calamansi, dalandan, and pomelo trees in mcdonald's farm is 8:3:2. There are a total of 793 trees in the said farm. How many trees of each kind are there?


5. If P89.96 is equivalent to $2, how much is $12 in pesos?


1
Expert's answer
2021-11-10T16:10:55-0500

(i) 3 men can work for 8 hours

1 man = 8*3 =24 hours

6 men = 24/6

=4hours.


(ii) Total amount divided=p5000

Ratio divided for Mina to Pina 3:5,

let Mina amount be 3y and Pina amount be 5y,

3y + 5y=8y

8y=5000

y=625

Mina =3(625) =p1875

Pina =5(625) =p3125


(iii) Rolando earning per hour =p75

Rolando earn 600 a day

total earn/total paid per hour

600/75=8.


(iv) Value of ratio x

 8x + 3x + 2x = 780

 

13x =780

 

x = 60

 

8(60) = 480 calamansi trees

3(60)=180 dalandan trees

 

There are 120 pomelo trees (60)

 

2(60) = 120 pomelo trees.


(v)  

P539.76

 

Step-by-step explanation:

 

P89.96/$2=P44.9/$1

 

This means $1=P44.98.

 

Therefore, in $12:

 

12(P44.98)=P539.76